"The fraternities and sororities of MSU face declining membership and want to know if the incoming freshman class will be a good recruiting ground. Not having enough money to survey all 1,600 incoming first year students, they commission you to survey a random sample of 150. You find that 35 of the 150 respondents on your survey are interested in social clubs."

What is your estimate of the number of first year students who would be interested in social clubs? (HINT: The upper and lower limits of your confidence interval are proportions. Convert proportions into numbers by multiplying with N = 1,600.)
Answer #1

Point estimate = 35/150 = 0.233

The 1st year total number of students who are interested in social clubs is estimated by = 1600 × 0.233 = 373.33 or 374.